Hey can someone help me please!
If the given angle is in standard position find two coterminal angles and two negative angles .
1.)-30 degrees
2.)5 pie/6
3.)-pie/4

Answer by lwsshak3(11628)

If the given angle is in standard position find two coterminal angles and two negative angles .
1.)-30 degrees
clockwise 360º+30º=-390º
clockwise 720º+30º=-750º
counter-clockwise 330º
counter-clockwise 360º+330º=690º
..
2.)5 pie/6
clockwise -7π/6
clockwise 12π/6+7π/6=-19π/6
counter-clockwise 12π/6+5π/6=17π/6
counter-clockwise 24π/6+5π/6=29π/6
..
3.)-pie/4
clockwise 8π/4+π/4=-9π/4
clockwise 16π/4+π/4=-17π/4
counter-clockwise 7π/4
counter-clockwise 8π/4+7π/4=15π/4